現代英語研究 IV 2 単位 (選択) 2 年 (後期)

現代英語研究 IV  スタージドナルド · 講師/ 人間文化学科

【授業目的】 This course introduces historical and contemporary features of foreign cultures so that students can compare and contrast foreign attitudes and Japanese attitudes. Students are encouraged to express their ideas in English orally and in writing. Formal and informal oral expression exercises are designed to make students comfrotable speaking in front of a group.

【授業概要】Class themes include: international attitudes to education, work,

morality, human rights, gender equity, politics. 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修上の注意】 A considerble number of the marks given in this course is for

participation in class. Good attendance is essential. 【到達目標】Upon successful completion of this course, students will have

experienced a course given entirely in English, learned significant historical and contemporary features of various foreign cultures, and compared conditions in Japan with conditions elsewhere. They will also have read authentic English language texts, led discussion in a seminar setting and practiced responding to queries regarding their experiences and opinions.

【授業計画】 1. A series of in-class oral presentation activities provides a basic structure.

2. Depending on class size, ther may also be lectures, note-taking exercises,

3. listening comprehension excercises and videos. 【成績評価】Participation (50%) Class Presentation (30%) Oral/Written Exams

(20%) 【再試験】 No 【教科書】 Class materials come primarily from students. There are some handouts

from the instructor. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218584 【連絡先】

⇒ スタージ (総合科学部 1 号館 2303, 088-656-7134, dws@ias.tokushima-u.a c.jp) ���� (オフィスアワー: 水曜日 16:30-17:30 または応相談)

現代英語研究 IV 2 units (selection) 2nd-year(2nd semester)

Donald Sturge · Associate Professor / Department of Human Sciences

Target⟩ This course introduces historical and contemporary features of foreign cultures so that students can compare and contrast foreign attitudes and Japanese attitudes. Students are encouraged to express their ideas in English orally and in writing. Formal and informal oral expression exercises are designed to make students comfrotable speaking in front of a group.

Outline⟩ Class themes include: international attitudes to education, work, morality,

human rights, gender equity, politics. �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Notice⟩ A considerble number of the marks given in this course is for participation

in class. Good attendance is essential. Goal⟩ Upon successful completion of this course, students will have experienced a

course given entirely in English, learned significant historical and contemporary features of various foreign cultures, and compared conditions in Japan with conditions elsewhere. They will also have read authentic English language texts, led discussion in a seminar setting and practiced responding to queries regarding their experiences and opinions.

Schedule⟩

1. A series of in-class oral presentation activities provides a basic structure.

2. Depending on class size, ther may also be lectures, note-taking exercises,

3. listening comprehension excercises and videos. Evaluation Criteria⟩ Participation (50%) Class Presentation (30%) Oral/Written

Exams (20%) Re-evaluation⟩ No Textbook⟩ Class materials come primarily from students. There are some handouts

from the instructor. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218584 Contact⟩

⇒ Sturge (Bldg. #1 2303, +81-88-656-7134, dws@ias.tokushima-u.ac.jp) ���� (Office Hour: Wednesday from 16:30-17:30 or by appointment)